# Machine-generated text detection prevents language model collapse

### v5

@@ 6.3 Results on collapse prevention
As explained in section 5, we assess our approach by adopting the decoding strategy that caused the most significant model collapse, i.e. pure sampling for GPT-2 and temperature sampling for SmolLM2, and for a more direct comparison, we also conduct experiments using top-$k$ decoding. At each generation $i$, we compare against the baseline of training on all samples in the pool of data $\mathcal{D}^{i}$ (Eq. 3). We also provide an “Oracle” performance, which represents a perfect machine-generated text detector that filters all synthetic samples.
@@
We evaluate recursive training in the partially synthetic setting under $3$ mixing settings ($\alpha\!=\!1$, $\beta\!=\!1$, $\gamma\!=\!0$), ($\alpha\!=\!0.5$, $\beta\!=\!1$, $\gamma\!=\!0$), ($\alpha\!=\!0.5$, $\beta\!=\!0.5$, $\gamma\!=\!0.5$). The task performance, data quality, and detector accuracy metrics over $10$ generations of partially synthetic training are depicted in Figure 2 for the scenario where human and synthetic samples have equal proportion (Appendices S4 and S6 contain results for the other $2$ scenarios). Weighted sampling (section 6.2) prevents model collapse and preserves the readability and diversity of the synthetic outputs, while the baseline degrades in task performance and data quality. The baseline generations become increasingly detectable, indicating divergence from human text. Notably, our method improves performance compared to training exclusively on human data (Oracle) across models and decoding strategies. These outcomes demonstrate both the value of using synthetic data in LLM training, but also the importance of selecting the right synthetic samples.
[…]
GPT-2 top-$k$5.5.5.5, $1$, 00 ${\color[rgb]{0,0,1}-7.28\%}$ ${\color[rgb]{0,0,1}+2.37\%}$ ${\color[rgb]{0,0,1}+4.54\%}$ ${\color[rgb]{0,0,1}-1.77\%}$ ${\color[rgb]{0,0,1}+0.46\%}$ ${\color[rgb]{0,0,1}+9.03\%}$
.5.5.5.5, .5.5.5.5, .5.5.5.5 ${\color[rgb]{0,0,1}-5.94\%}$ ${\color[rgb]{0,0,1}+1.72\%}$ ${\color[rgb]{0,0,1}+2.43\%}$ ${\color[rgb]{0,0,1}-4.28\%}$ ${\color[rgb]{0,0,1}+3.99\%}$ ${\color[rgb]{0,0,1}+10.16\%}$
$1$, $1$, 00 ${\color[rgb]{0,0,1}-4.45\%}$ ${\color[rgb]{0,0,1}+1.49\%}$ ${\color[rgb]{0,0,1}+3.59\%}$ ${\color[rgb]{0,0,1}-3.58\%}$ ${\color[rgb]{0,0,1}+1.36\%}$ ${\color[rgb]{0,0,1}+6.76\%}$
pure sampling .5.5.5.5, $1$, 00 ${\color[rgb]{0,0,1}-7.41\%}$ ${\color[rgb]{0,0,1}+1.50\%}$ ${\color[rgb]{1,0,0}-1.30\%}$ ${\color[rgb]{1,0,0}+36.71\%}$ ${\color[rgb]{0,0,1}+74.06\%}$ ${\color[rgb]{0,0,1}+50.23\%}$
.5.5.5.5, .5.5.5.5, .5.5.5.5 ${\color[rgb]{0,0,1}-6.54\%}$ ${\color[rgb]{0,0,1}+1.50\%}$ ${\color[rgb]{1,0,0}-1.07\%}$ ${\color[rgb]{1,0,0}+21.39\%}$ ${\color[rgb]{0,0,1}+16.38\%}$ ${\color[rgb]{0,0,1}+25.45\%}$
$1$, $1$, 00 ${\color[rgb]{0,0,1}-25.65\%}$ ${\color[rgb]{0,0,1}+0.96\%}$ ${\color[rgb]{1,0,0}-0.71\%}$ ${\color[rgb]{1,0,0}+20.70\%}$ ${\color[rgb]{0,0,1}+16.42\%}$ ${\color[rgb]{0,0,1}+20.99\%}$
SmolLM2 $360$M top-$k$5.5.5.5, $1$, 00 -4.60% +0.68% +7.06% -1.17% +15.28% -0.73%
.5.5.5.5, .5.5.5.5, .$5$4.37% +0.62% +4.25% -0.42% +2.05% -0.74%
$1$, $1$, 00 -3.05% +0.54% +6.20% -1.74% +10.62% -0.61%
temperature .5.5.5.5, $1$, 00 -3.91% +0.55% +14.62% -1.46% +20.92% -1.86%
.5.5.5.5, .5.5.5.5, .$5$3.24% +0.48% +7.44% -1.98% -0.33% -1.35%
$1$, $1$, 00 -2.23% +0.39% +8.55% -1.80% +13.98% -0.34%
SmolLM2 $135$M top-$k$ $1$, $1$, 00 -4.19% +0.75% +9.96% -2.15% +13.08% -0.76%
SmolLM2 $1.7$B top-$k$ $1$, $1$, 00 -0.85% +0.06% +3.44% -1.11% +8.46% -0.65%
Table 3: Percentage of change in data quality when using our proposed mitigation strategy versus the baseline. Results are shown for top-$k$ decoding and pure sampling/temperature for different values of $\alpha$, $\beta$, and $\gamma$ (blue / red: positive / negative results, $\uparrow$ / $\downarrow$: higher / lower is better).
@@
Table 3 enumerates the percentage difference for the final models under the baseline strategy vs. using our approach. We also enumerate the exact scores in Table S2. Our method improves the data quality and model performance across all metrics, except for pure sampling with GPT-2, where the baseline shows higher diversity and Self-BLEU but at the cost of lower MAUVE, readability, and task performance, indicating degraded quality. Notably, we observe that mixing cross-generational data has a minimal effect on the extent of model collapse compared to training solely on the previous generation, contrasting with the findings of Gerstgrasser et al. (2024). However, we note that Gerstgrasser et al. (2024) did not constrain the sample size. Our experiments adopt a more realistic and less extreme setting by sampling a fixed dataset under different mixing scenarios.
Additionally, we evaluate the effectiveness of our method at different model scales of SmolLM2 ($135$M, $360$M, $1.7$B) for a fixed decoding strategy and data mixing setting. The results are enumerated at the bottom of Table 3 and the performance across generations is depicted in Figure S5. Our method prevents model collapse across all SmolLM variants. Notably, we find that smaller models exhibit greater relative improvements using our method compared to both the Oracle and the baseline.
@@
Figure 3: Perplexity distribution for machine-generated text of GPT-2 at generations 00, $1$, and $9$ in the partially synthetic scenario $\left(\alpha\!=\!0.5,\beta\!=\!1,\gamma\!=\!0\right)$. Perplexity is evaluated using the model trained on human text ($p^{0}$).
@@
In line with previous research Shumailov et al. (2024), we also study the perplexity distribution of the synthetic data at each generation using model $p^{0}$ that was trained on human data. Figure 3 depicts these distributions for generations 00, $1$, and $9$ for the GPT-2 model with top-$k$ and pure sampling, compared to the baseline (Appendix S3 contains the results for SmolLM2). For top-$k$ decoding, similarly to Shumailov et al. (2024), we observe that for the baseline, perplexity shifts towards regions of lower scores and the distribution becomes more peaked, showing signs of early model collapse. For pure sampling, on the other hand, we observe that the baseline distribution shifts to higher perplexity scores and displays increased variance. This is an interesting finding that demonstrates that by removing truncation from the decoding strategy, the narrowing effect of model collapse is diminished, and instead, model collapse is reflected by long-tail incoherent text that is completely distinct from the original human samples. By deploying our mitigation strategy, however, we observe very little change in the perplexity distribution for both sampling strategies.
@@ 7 Conclusion
This work investigates model collapse across three dimensions: model performance, data quality, and resemblance to human samples. Through our analysis, we found that the extent of model collapse and the effect on the data distribution is influenced by the decoding strategy. Truncating can lead to peaked distributions and repetitive models while pure sampling can result in high perplexity and verbose outputs with low resemblance to human data. Using the decoding strategies that resulted in the most extreme collapse, we evaluated the partially synthetic scenario, where human data is mixed into the training data. We designed a novel method to mitigate model collapse based on resampling the training distribution using the predictions of a machine-generated text detector. We have validated our method on two popular model variants (GPT-2 and SmolLM2) across a range of decoding strategies and model sizes, showing that we can prevent model collapse in all cases. When there is an equal ratio of human to synthetic samples in the training pool, our method results in improved model performance compared to training only on the human data.
@@
Finally, our method depends on the performance of the machine-generated text detector used to estimate the importance weights. Any inaccuracies or biases in the detector’s predictions directly affect the quality of the resampling process, potentially leading to suboptimal mitigation of model collapse. Since detector performance varies with domain, LLM architecture, and decoding strategy, the generalisability of our approach is closely tied to the detector’s robustness. The primary focus of this work, however, is not on optimising the detector itself, but on demonstrating that detector-based resampling can effectively mitigate model collapse. Future work could investigate improved methods of detection, such as using adaptive or ensemble-based detectors to improve reliability across data regimes and LLMs.
Ethics Statement
The authors declare no competing interests. The Wikitext-2 dataset Merity et al. (2016) used for the experiments is licensed under the Creative Commons Attribution-ShareAlike License (CC BY-SA 4.0). GPT-2 is licensed under the MIT license and SmolLM2 is licensed under APACHE 2.0. AI assistants were not used for this research.
@@ A.1 Dataset
All our model collapse experiments use the raw variant of the WikiText-2 dataset Merity et al. (2016). We train models on the training set, consisting of $36{,}718$ documents, and evaluate on the test set of $4{,}358$ documents. The WikiText-2 dataset was extracted from the ‘Good’ or ‘Featured’ article criteria specified by editors on Wikipedia and only covers the English language.
@@ A.2 LLMs
GPT-2 (Generative Pre-trained Transformer 2) Radford et al. (2019) is a decoder-only transformer-based language model. GPT-2 demonstrated that large-scale language models could perform various language tasks without task-specific training. We use the base variant, which contains $124$M parameters. SmolLM2 Allal et al. (2025) is a family of compact and efficient language models developed by Hugging Face, available in three sizes: $135$M, $360$M, and $1.7$B parameters. The majority of our experiments use the $360$M parameter variant unless specified otherwise.
@@ A.3 Hyperparameters
In our experiments, we conduct full fine-tuning using a learning rate of $5\!\times\!10^{-5}$, batch size of $8$ and a dropout rate of $0.1$. For the AdamW optimizer, we set $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. Each model was trained for $1$ epoch with the hyperparameters fixed for all experiments. We conducted $10$ iterations of recursive training.
@@ B.1 Pre-trained models
Robustly Optimized BERT pre-training Approach (RoBERTa) by Liu et al. (2019) improves on the pre-training phase of BERT Devlin et al. (2019), an encoder-only transformer model that leverages masked language models to enable pre-trained deep bidirectional representations. The RoBERTa model optimised the pre-training procedure for BERT by training the model for longer and on more data, changing the masking pattern, and removing the next sentence prediction objective. We use the base variant which has $125$M parameters.
Decoding-enhanced BERT with Disentangled Attention (DeBERTav3) by He et al. (2023), is a BERT-based encoder only model enhanced with disentangled attention. DeBERTav3 model improves on DeBERTa by using Enhanced Mask Decoding and an ELECTRA-style pre-training objective, Replaced Token Detection, instead of Masked Language Modelling. We use the base variant which contains $86$M backbone parameters with an embedding layer of $98$M parameters.
ModernBERT Warner et al. (2024) is a recent addition to the encoder-only transformer models that has been designed to increase downstream performance and efficiency on GPUs, particularly in long-context scenarios due to its $8{,}192$ native sequence length. The model was trained on $2$ trillion tokens and improves on the original BERT architecture with rotary positional embeddings (RoPE), unpadding, GeGLU layers and alternating local-global attention demonstrating SOTA performance amongst encoder models across a range of classification and retrieval tasks. We conduct experiments with the base variant, which contains $150$M parameters.
@@ B.2 Hyperparameters
Each model was fine-tuned for $5$ epochs. We select the best model based on the highest AUC on the validation set. Optimisation was performed using AdamW by setting $\beta_{1}=0.9$, $\beta_{2}=0.98$, $\epsilon=10^{-6}$, and the weight decay to $10^{-2}$. These parameters were chosen based on prior work Warner et al. (2024). The label smoothing parameter $\alpha$ was set to 0.1, the seed was fixed at $42$ and the training batch size to $16$. The learning rate was set based on a hyperparameter sweep over [$1$, $1.5$, $2$, $3$, $4$]×$\times 10^{-5}$. For ModernBERT the best-performing learning rate was $10^{-5}$. We implemented temperature scaling by learning the temperature parameter using L-BFGS optimisation on the validation set. This was run for $50$ iterations with a learning rate of $0.01$.
@@ B.3 Dataset
We trained and evaluated the machine-generated text detectors on the MAGE dataset Li et al. (2024), which is based on documents from $10$ domains: opinion statements (CMV & Yelp reviews dataset), news articles (XSum & TLDR dataset), question answering (ELI5), story generation (Reddit WritingPrompts & ROC), commonsense reasoning (HellaSwag), knowledge illustration (SQuAD) and Scientific writing (SciGen). The authors sampled $1{,}000$ texts from each domain (apart from opinion statements and news articles with $804$ and $777$ samples respectively) and generated text using $27$ LLMs from $7$ model families, which include OpenAI, LLaMA, GLM, FLAN-T5, OPT, BigScience and EleutherAI. For each human-written sample in the dataset, they generate a machine-generated version by providing the first $30$ tokens of human-written text as context to the LLM. In addition, for the OpenAI models, they implemented two other prompt strategies for relevant domains: ‘topical’ prompts such as an argument or news title and ‘specified’ prompts which contain information about the domain source. This results in $33{,}000$ ($=27{,}000+3\!\times\!2\!\times\!1{,}000$) machine-generated samples per source before processing and filtering. The authors split the dataset into train, validation and test splits in the ratio $80$:$10$:$10$. To mitigate data imbalance in the validation and test sets they sample additional human data from each data source. The resulting test set contains $28{,}741$ human and $28{,}078$ machine-generated samples ($49\%$ machine-generated). The training set, however, is $71\%$ machine-generated. The total dataset consists of $154{,}078$ human-written and $294{,}381$ machine-generated texts. In addition to the previously described test set, we also evaluate our detector on their more challenging test set containing text from four unseen domains (CNN/DailyMail, DialogSum, PubMedQA, IMDB) and generated by an unseen model (GPT-4). This out-of-distribution test set contains $762$ human and $800$ machine-generated samples.
@@
When evaluating the ModernBERT model fine-tuned on MAGE on the SmolLM2 models, we observed a drop in performance compared to GPT-2, with large variability across decoding strategies and model size. For SmolLM2 $360$M the detector achieved a classification accuracy of .601.601.601.601 for top-$k$ decoding and .399.399.399.399 for temperature sampling. To ameliorate this, we finetuned a new ModernBERT model on a larger corpus, containing the MAGE dataset and a subset of the RAID dataset Dugan et al. (2024) for the SmolLM2 models. The RAID dataset is the largest machine-generated text detection dataset, includes text samples generated by $11$ LLMs with $4$ decoding strategies, and spans text across $8$ domains. Additionally, RAID includes $11$ types of adversarial attacks, such as homoglyph substitutions, number insertions, article deletions, and paraphrasing. We partitioned the dataset into training, validation, and test splits in the ratio $80$:$10$:$10$, ensuring no cross-contamination of text segments generated from the same source of text across splits. We balanced each split so that it contained an equal number of human and machine samples, stratified across model, decoding strategy, source domain, and adversarial attack (the whitespace and paragraph attacks were included). This resulted in balanced train, validation, and test splits comprising $128{,}352$, $16{,}044$, and $16{,}056$ samples, respectively.

### v6

@@ 6.3 Results on collapse prevention
As explained in section 5, we assess our approach by adopting the decoding strategy that caused the most significant model collapse, i.e. pure sampling for GPT-2 and temperature sampling for SmolLM2. For a more direct comparison, we also conduct experiments using top-$k$ decoding for both models. At each generation $i$, we compare against the baseline of training on all samples in the pool of data $\mathcal{D}^{i}$ (Eq. 3). We also provide an ‘Oracle’ performance, which internally deploys an infallible machine-generated text detector that filters all synthetic samples to ensure that the training dataset only contains human-written texts.
@@
We evaluate recursive training in the partially synthetic setting under $3$ mixing settings: $\left(\alpha\!=\!1,\beta\!=\!1,\gamma\!=\!0\right)$, $\left(\alpha\!=\!0.5,\beta\!=\!1,\gamma\!=\!0\right)$, and $\left(\alpha\!=\!0.5,\beta\!=\!0.5,\gamma\!=\!0.5\right)$. The task performance, data quality, and detector accuracy metrics over $10$ generations of partially synthetic training are depicted in Figure 3 for the scenario where human and synthetic samples have equal proportion (Appendices S3 and S4 contain results for the other two scenarios). The results show that our method of weighted resampling (section 6.2) prevents model collapse and preserves the readability and diversity of the synthetic outputs, while the baseline degrades in task performance and data quality. Over the generations, the outputs in the baseline become increasingly detectable, indicating a divergence from the characteristics of human-written text. In addition, our method improves performance compared to training exclusively on the human samples (‘Oracle’), demonstrating both the value of using synthetic data in LLM training, but also the importance of selecting the right synthetic samples.
@@
Table 3 enumerates the percentage of difference across various performance metrics for the final model generations under the baseline strategy vs. using our approach. We also enumerate the corresponding raw values in Table S2. Our method improves the data quality and model performance across all metrics, except for pure sampling with GPT-2, where the baseline shows higher diversity and Self-BLEU but at the cost of lower MAUVE, readability, and task performance, indicating degraded model quality. Notably, in the setting where data is accumulated across generations $\left(\alpha\!=\!0.5,\beta\!=\!0.5,\gamma\!=\!0.5\right)$, we observe that mixing cross-generational data has a minimal effect on the extent of model collapse compared to training solely on the previous generation. This contrasts with the findings of Gerstgrasser et al. (2024), who suggest that accumulating all synthetic samples alongside the original human samples can avoid model collapse. However, we note that Gerstgrasser et al. (2024) did not constrain the sample size and trained on increasing volumes of data at each generation. Our experiments adopt a more realistic setting by sampling a fixed dataset size for each generation under different mixing scenarios.
GPT-2 SmolLM2
Baseline
Refer to caption Refer to caption Refer to caption Refer to caption
Mitigation
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 4: Perplexity distribution of the machine-generated text at generations 0, 11, and $9$, evaluated using the model trained on human text ($p^{0}$). Results are shown for top-$k$, and pure sampling/temperature decoding for GPT-2/SmolLM2 under the baseline strategy (top) and our importance resampling method (bottom) for the partially synthetic setting $\left(\alpha\!=\!0.5,\beta\!=\!1,\gamma\!=\!0\right)$.
Additionally, we evaluate the effectiveness of our method at different model scales of SmolLM2 (135135M, 360360M, 1.71.7B) using top-$k$ decoding due to its consistent effect across models and with a fixed data mixing setting $\left(\alpha\!=\!1,\beta\!=\!1,\gamma\!=\!0\right)$. The results are enumerated at the bottom of Table 3 and the performance across generations is depicted in Figure S5. Our method prevents model collapse across all SmolLM model sizes, improving perplexity, accuracy, diversity, Self-BLEU, and MAUVE, with only minor reductions in readability. Notably, we find that smaller models exhibit greater relative improvements using our method compared to both the ‘Oracle’ and the baseline, with the 135135M variant reducing perplexity by $4.19\%$ and increasing diversity by $9.96\%$. We hypothesise that this is because the data generated by smaller models is inherently of lower quality, leading to larger gains over the baseline by resampling the training data to reduce the proportion of synthetic samples.
@@
In line with previous research Shumailov et al. (2024), we also study the perplexity distribution of the synthetic data at each generation as evaluated by the original model $p^{0}$ trained on the human data. Figure 4 depicts these distributions for generations 0, 11, and $9$ for GPT-2 (top-$k$ and pure sampling decoding) and SmolLM2 (top-$k$ and temperature sampling decoding) compared to the baseline. For top-$k$ decoding and temperature sampling, similarly to Shumailov et al. (2024), we observe that over the generations, models produce samples that the original model, $p^{0}$, trained on the human data, would also be more likely to produce. This is depicted through perplexity scores, which shift towards regions of lower values, while their distribution becomes more peaked, showing signs of early model collapse. For pure sampling, on the other hand, we observe that the distribution shifts to higher perplexity scores and displays increased variance. This is an interesting finding that demonstrates that by removing truncation or temperature from the decoding strategy, the narrowing effect of model collapse is diminished, and instead, model collapse is reflected by long-tail incoherent text that is completely distinct from the original human samples. By deploying our mitigation strategy, however, we observe very little change in the perplexity distribution for all three sampling strategies.
@@ 7 Conclusion
This work investigates model collapse across three dimensions: model performance, data quality, and semantic similarity to human samples. Our analysis highlights that the extent of model collapse and the effect on the output distribution are influenced by the decoding strategy. Truncating can lead to peaked distributions and repetitive models, while pure sampling can result in high perplexity and verbose outputs with low resemblance to human text. Using the decoding strategies that resulted in the most extreme collapse, we evaluated the partially synthetic scenario, where human and LLM samples are mixed into the training set. We proposed a method to mitigate model collapse by resampling the training distribution with importance weights guided by the predictions of a machine-generated text detector. We validated our method on two language model families (GPT-2 and SmolLM2) across a range of decoding strategies and model sizes (124124M to 1.71.7B), showing that model collapse can be prevented in all cases. Additionally, when the ratio of human to synthetic samples in the initial training pool is equal, our resampling method yields an improved performance compared to training exclusively on human data.
@@
Lastly, as our model collapse experiments require full fine-tuning of a new model at each generation (for all $10$ generations), it becomes computationally expensive and time-intensive to evaluate large-scale models. On top of that, for the experiments we present in our paper, computational load increases significantly because we evaluate $6$ decoding strategies and $4$ data mixing settings. For this reason, the largest model evaluated is under 22B parameters, which is at the limit of our available resources. Future work could extend this analysis to larger models to validate the generalisability of these findings with large-scale SOTA models.
Acknowledgments
G. Drayson is funded by the EPSRC grant “AI Centre for Doctoral Training in Foundational Artificial Intelligence” (EP/S021566/1). The authors would like to thank the area chairs who reviewed this paper for their constructive feedback.
@@ A.1 Dataset
All our model collapse experiments use the raw variant of the WikiText-2 dataset Merity et al. (2017), which is licensed under the Creative Commons Attribution Share Alike Licence (CC BY-SA 4.0). We train models on the training set, consisting of 36,71836{,}718 documents, and evaluate on the test set of 4,3584{,}358 documents. The WikiText-2 dataset was extracted from the ‘Good’ or ‘Featured’ article criteria specified by editors on Wikipedia and only covers the English language.
@@ A.2 LLMs
GPT-2 (Generative Pre-trained Transformer 2; Radford et al. 2019) is a decoder-only transformer-based language model that is available under the MIT licence. GPT-2 demonstrated that large-scale language models could perform various language tasks without task-specific training. We use the base variant, which contains 124124M parameters. SmolLM2 Allal et al. (2025) is a family of compact and efficient language models developed by Hugging Face, available in three sizes: 135135M, 360360M, and 1.71.7B parameters and is available under the APACHE 2.0 licence. The majority of our experiments use the 360360M parameter variant unless specified otherwise.
@@ A.3 Hyperparameters
In our experiments, we conduct full fine-tuning using a learning rate of $5\!\times\!10^{-5}$, batch size of $8$ and a dropout rate of $0.1$. For the AdamW optimizer, we set $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. Each model is trained for $1$ epoch with the hyperparameters fixed for all experiments. We conducted $10$ iterations of recursive training.
@@ B.1 Pre-trained models
Robustly Optimized BERT pre-training approach (RoBERTaLiu et al. 2019) improves on the pre-training phase of BERT Devlin et al. (2019), an encoder-only transformer model that leverages masked language models to enable pre-trained deep bidirectional representations. The RoBERTa model optimised the pre-training procedure for BERT by training the model for longer and on more data, changing the masking pattern, and removing the next sentence prediction objective. The model is licensed under the MIT licence. We use the base variant which has 125125M parameters.
Decoding-enhanced BERT with Disentangled Attention (DeBERTav3He et al. 2023) is a BERT-based encoder only model enhanced with disentangled attention. DeBERTav3 improves on DeBERTa by using Enhanced Mask Decoding and an ELECTRA-style pre-training objective, Replaced Token Detection, instead of Masked Language Modelling. The model is licensed under the MIT licence. We use the base variant which contains 8686M backbone parameters with an embedding layer of 9898M parameters.
ModernBERT Warner et al. (2024) is a recent addition to the encoder-only transformer models that has been designed to increase downstream performance and efficiency on GPUs, particularly in long-context scenarios due to its 8,1928{,}192 native sequence length. The model was trained on $2$ trillion tokens and improves on the original BERT architecture with rotary positional embeddings (RoPE), unpadding, GeGLU layers, and alternating local-global attention demonstrating SOTA performance amongst encoder models across a range of classification and retrieval tasks. The model is available under the APACHE 2.0 licence. We conduct experiments with the base variant, which contains 150150M parameters.
@@ B.2 Hyperparameters
[…]
@@ B.3 Dataset
We trained and evaluated the machine-generated text detectors on the MAGE dataset Li et al. (2024), which is based on documents from $10$ domains: opinion statements (CMV & Yelp reviews dataset), news articles (XSum & TLDR dataset), question answering (ELI5), story generation (Reddit WritingPrompts & ROC), commonsense reasoning (HellaSwag), knowledge illustration (SQuAD), and Scientific writing (SciGen). The authors sampled 1,0001{,}000 texts from each domain (apart from opinion statements and news articles with $804$ and $777$ samples respectively) and generated text using $27$ LLMs from $7$ model families, which include OpenAI, LLaMA, GLM, FLAN-T5, OPT, BigScience, and EleutherAI and their dataset is available under the APACHE 2.0 licence. For each human-written sample in the dataset, they generate a machine-generated version by providing the first $30$ tokens of human-written text as context to the LLM. In addition, for the OpenAI models, they implemented two other prompt strategies for relevant domains: ‘topical’ prompts such as an argument or news title and ‘specified’ prompts which contain information about the domain source. This results in 33,00033{,}000 ($27{,}000+3\!\times\!2\!\times\!1{,}000$) machine-generated samples per source before processing and filtering. The authors split the dataset into train, validation and test splits in the ratio 8080:1010:$10$. To mitigate data imbalance in the validation and test sets they sample additional human data from each data source. The resulting test set contains 28,74128{,}741 human, and 28,07828{,}078 machine-generated samples ($49\%$ machine-generated). The training set, however, is $71\%$ machine-generated. The total dataset consists of 154,078154{,}078 human-written and 294,381294{,}381 machine-generated texts. In addition to the previously described test set, we also evaluate our detector on their more challenging test set containing text from four unseen domains (CNN/DailyMail, DialogSum, PubMedQA, IMDB) with machine-generated text from an unseen model (GPT-4). This out-of-distribution test set contains $762$ human and $800$ machine-generated samples.
@@
When evaluating the ModernBERT model fine-tuned on MAGE on the SmolLM2 models, we observed a drop in detection performance compared to GPT-2 text, with large variability across decoding strategies and model size. For SmolLM2 360360M the detector achieved a classification accuracy of .601.601 for top-$k$ decoding while for temperature sampling the accuracy was .399.399. To ameliorate this, we finetuned a new ModernBERT model on a larger corpus, containing the MAGE dataset and a subset of the RAID dataset Dugan et al. (2024) for the SmolLM2 models. The RAID dataset is the largest machine-generated text detection dataset including text samples generated by $11$ LLMs with $4$ decoding strategies, and spans text across $8$ domains. Additionally, RAID includes $11$ types of adversarial attacks, such as homoglyph substitutions, number insertions, article deletions, and paraphrasing and is licensed under the MIT licence. We partitioned the dataset into training, validation, and test splits in the ratio 8080:1010:$10$, ensuring no cross-contamination of text segments generated from the same source document across splits. We balanced each split so that it contained an equal number of human and machine samples, stratified across model, decoding strategy, source domain, and adversarial attack (the whitespace and paragraph attacks were included). This resulted in balanced train, validation, and test splits comprising 128,352128{,}352, 16,04416{,}044, and 16,05616{,}056 samples, respectively.